A random variable has a triangular probability density function with a = 50, b = 375, and m = 250. What is the probability that the random variable will assume a value between 100 and 300?
Answer #1

from above

P(X<100 ) =(x-a)2/((b-a)*(m-a)) = 0.0385
P(X<300)=1-(b-x)2/((b-a)*(b-m))= 0.8615

P(100<X<300) =P(X<300)-P(X<100) =0.8615-0.0385 =0.8230